OpenStudy (anonymous):

If f(x)=x4-14x3+72x2+5x-8, find the largest interval on which f(x) is concave downward. If we write the interval as (a,b), then a= ? and b=?

OpenStudy (anonymous):

take derivative twice and find where the equation is negative between two zeros

OpenStudy (anonymous):

Okay, I got the 2nd der. to be [12x ^{2}-86x=-144\]...what do i do next?

OpenStudy (amistre64):

set it to zero

OpenStudy (amistre64):

i always have to recall, i knw x^2 is cave up, and its 2nd derivative is positive; so if 2nd derivative is positive you got a cave up, negative is a cave down

OpenStudy (amistre64):

where 2nd D is 0 you got inflections

OpenStudy (amistre64):

most likely got inflections but could have some false readings that need to be dbl chkd

OpenStudy (anonymous):

I'm not sure that I'm following what you're saying...

myininaya (myininaya):

\[f'(x)=4x^3-42x^2+154x+5\] \[f''(x)=12x^2-84x+154\] set f''=0 \[12x^2-84x+154=0\] \[6x^2-42x+77-0\] \[x=\frac{-(-42)-\sqrt{(-42)^2-4(6)(77)}}{2(6)}\]

OpenStudy (anonymous):

your second derivative is a quadratic function that faces up. as such it will be negative between the zeros and positive outside them

OpenStudy (anonymous):

so your original function will first be concave up ( a stupid term, because it sounds just like "concave down", so what is the word "concave" doing there) then concave down (frowning, leaning right, spilling water) then concave up again.

OpenStudy (anonymous):

got it! thanks!